The displacement x of a particle in time t is given by 10t2-4t -x =0 where x is i metre and t in second the displacement covered by the body in 4th second of motion

Grade:11

2 Answers

Bhaskar mishra
4 years ago
Since, t=4second given so
Put the value of t=4in the given equation and find that 
The value of x=144
10*4*4-4*4-x=0
160-16-x=0
X=160-16
X=144
Bhaskar mishra
4 years ago
We know that the for 4th second there we use formulae of s4th - s3rd and that is 
We put the value of t =4 in given equation and t=3in given equation and subtract the value of S3 from S4
Then
10*4*4-4*4-x=0
X=144 it is the value of S4
Now, 10*3*3-3*3-x=0
X=78
So, the final answer is 144-78=66